(1 point) Find a particular solution for each differential equation y + 3y = 5t, then yp = y - y = 3et sin(t), then yp = y
Answer #1

diff.eq. find a particular sol. for each 9+ 3y = 5t sol y+3y=5t - A.E.S m²+3m=o mcm+3)=0 m=o,3 Je = c +cest Now we find up bNow ① gluy= betsint , hou Up = ylly= 3e sint - 0 A.E. is m?i=0 m=1 ..m= I) = m = 1,-) You cieti cset we find Sp, by method of-2A-B=3 - 2(2B)-B=3 =) -53=3 =) B: - A=28=A = 2(0) = -6 - A = , B=-3 ->p = get cosct) - Besinct) -Yp - let-cascto a sinct))

~ y44y = cos(at), yp: solo y+4y=coscat) - A.E.18 m²+4=0 m²=-4 m=zzi .: 9c-4,28(at)+cą Sin (et) Now we find 9p:-) by method ocauating coefficients of similar teams on both sides s Sinat :-) -4A=0 ) A=0 2 coat :-) 4B=1 ) B= // A=0, B=4 Yp= 1 + Sin cat